# Tiny Time Mixers (TTMs): Fast Pretrained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

### 11 Jan 2024 v2

@@ 1 Introduction
Related Work: Recent advances in multivariate forecasting have been marked by the advent of transformer-based Vaswani et al. (2017) approaches, exemplified by models like PatchTST Nie et al. (2022), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and FEDFormer Zhou et al. (2022). These models have demonstrated notable improvements over traditional statistical and ML methods. Furthermore, architectures based on MLP-Mixer Tolstikhin et al. (2021), such as TSMixer Ekambaram et al. (2023), have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts. However, none of these advanced approaches have successfully demonstrated the ability to create general pretrained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks. This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of pretraining TS data. There are existing self-supervised pretraining TS approaches using masked modeling and contrastive learning techniques such as SimMTM Dong et al. (2023) and TF-C Zhang et al. (2022) that offer transfer learning between two datasets when carefully selected based on the dataset properties. However, they fail to provide universal transfer learning capabilities across datasets. Consequently, there has been a recent growing trend to employ pretrained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task. These universal cross-transfer approaches, specifically recent works such as LLMTime Gruver et al. (2023) and GPT4TS Zhou et al. (2023) yield promising results in few/zero-shot forecasting approaches. These models are bootstrapped from GPT-2/3 or LLAMA-2 with suitable tokenization strategies to adapt to time-series domains.
@@
However, these LLM based TS approaches do not explicitly handle channel correlations and exogenous support in the context of multivariate forecasting. Moreover, these large models, with billions of parameters, demand significant computational resources and runtime. Hence, in this paper, we focus on building pretrained models from scratch solely using TS data. Unlike language, which has abundant public pretraining data in terabytes, time-series data is relatively scarce, very diverse and publicly limited. Its scarcity leads to overfitting when pretraining “large” models solely on time-series data. This prompts a question: Can smaller models pretrained purely on limited public diverse TS datasets give better zero/few-shot forecasting accuracy? Surprisingly, the answer is yes! Toward this, we propose Multi-level Tiny Time Mixers (TTM), a significantly smaller model ($\leq$1M parameters) based on the lightweight TSMixer architecture, exclusively trained on diverse TS corpora for effective zero/few-shot multivariate TS forecasting via transfer learning.
@@
In particular, TTM is pretrained using multiple public datasets ($\sim$ 240M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ Godahewa et al. (2021)). Note that the datasets exhibit considerable diversity in terms of characteristics, such as the different domains, temporal resolution [^3] Resolution refers to the sampling rate of the input time series (e.g., hourly, 10 minutes, 15 minutes, etc.) (spanning from second to daily), lengths, and number of channels. Pretraining on such heterogeneous datasets cannot be handled directly by TSMixer or existing state-of-the-art (SOTA) models. Hence, TTM proposes the following enhancements to the TSMixer architecture: (i) Adaptive Patching across layers, considering the varied suitability of patch lengths for different datasets, (ii) Dataset Augmentation via Downsampling to increase coverage and samples across different resolutions, (iii) Resolution Prefix Tuning to explicitly embed resolution information in the first patch, facilitating resolution-conditioned modeling, particularly beneficial in scenarios with short history lengths. Moreover, our approach leverages multi-level modeling, where TTMs are first pretrained in a channel-independent way and then seamlessly integrate channel mixing during fine-tuning to model target data-specific channel-correlations and exogenous infusion
@@
  • Amidst the prevalence of large pretrained models demanding significant compute and training time, our work is the first to showcase the efficacy of building Fast and Tiny Pretrained models ($\leq$1M parameters) exclusively trained on Public TS datasets in a flash of just 4-8 hours (using 6 A100 GPUs). TTM successfully demonstrates transfer learning to diverse, unseen target datasets for zero/few-shot forecasting, addressing the data scarcity issues prevalent in time series.
  • Pretraining on heterogeneous multi-resolution datasets cannot be handled effectively by TSMixer or other SOTA models. Hence, we propose various architectural and training enhancements, such as adaptive patching, data augmentation via downsampling, and (an optional) resolution prefix tuning for robust pretraining.
@@ 2 TTM Components
Let $\bm{X}\in\mathbb{R}^{c\times sl}$ be a multivariate time series of length $sl$ and $c$ number of channels. The forecasting task can be formally defined as predicting the future values $\bm{Y}\in\mathbb{R}^{c^{\prime}\times fl}$ given the history $\bm{X}$. Here, $fl$ denotes the forecast horizon and $c^{\prime}$ denotes the forecast channels, where $c^{\prime}\leq c$. The predictions from the model are denoted by $\hat{\bm{Y}}\in\mathbb{R}^{c^{\prime}\times fl}$. In a general multivariate forecasting task, each channel or variable falls into one of the following categories: (a) Target variables (mandatory):corresponding to the channels ($c^{\prime}$) for which forecasts are required, (b) Exogenous variables (optional):encompassing (i) uncontrolled variables that may influence the forecasts and assumed to be known or estimated for the forecast period (e.g. weather), and (ii) control variables whose future values during the forecast horizon can be manipulated to govern the behavior of the target variables. (e.g. discount in sales forecasting, operator controls in industrial applications). In TTM, uncontrolled and control variables are treated similarly, as both are considered available during forecasting.
@@ 3 TTM Workflows
TTM works in two stages: pretrain and finetune (Figure 1).
3.1 Pretraining Workflow
@@
In the pretraining stage, we train the model on a large collection of public datasets from the Monash data repositoryGodahewa et al. (2021). Since the primary focus of TTM is forecasting, pretraining is modeled with a direct forecasting objective. TTM is first pretrained in a univariate fashion with independent channels on all the existing datasets. Due to varied channel counts in pretraining datasets, modeling multivariate correlations is not feasible here; it is addressed later during fine-tuning. Multivariate pretraining datasets are initially transformed into independent univariate time series $(\bm{X}_{1},\cdots,\bm{X}_{N})\in\mathbb{R}^{c(=1)\times sl}$. These are pre-processed (Section 2.2), and subsequently fed into the TTM backbone for multi-resolution pretraining. The output of the backbone $\bm{X}_{h}^{L}\in\mathbb{R}^{(c=1)\times n\times hf}$ is passed through the decoder and forecast head to produce the forecast $\hat{\bm{Y}}\in\mathbb{R}^{(c=1)\times fl}$ which is then reverse-normalized to bring back to the original scale. We pretrain the TTM with mean squared error (MSE) loss function calculated over the forecast horizon: $\mathcal{L}=||\bm{Y}-\hat{\bm{Y}}||_{2}^{2}$. Thus for a given input context length $sl$ and forecast length $fl$, we get a pretrained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pretraining data.
3.1.1 Multi-Resolution Pretraining via TTM Backbone
@@
The majority of the pretraining happens in the TTM backbone. The primary challenge with the proposed pretraining technique is that the pretraining data is diverse and has multiple resolutions. There are two main options for pretraining: conducting separate pretraining for each resolution type or pretraining using all resolution data collectively. While it’s common to train a model per resolution type to overcome challenges in learning diverse seasonal patterns, this leads to diminished training data for each resolution due to limited data availability. Consequently, this motivated the exploration of pretraining a single model using datasets from all resolutions. To achieve this, we propose the following 3 enhancements.
@@
Adaptive Patching: The TTM backbone is crafted with an adaptive patching architecture where different layers of the backbone operate at varying patch lengths and numbers of patches. Since each dataset in the pretraining corpora may perform optimally at a specific patch length, this approach greatly aids in generalization when diverse datasets with different resolutions are introduced. As shown in Figure 1(b), the patched data $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ is passed through a embedding layer to project it to the patch hidden dimension, $\bm{X}_{h}\in\mathbb{R}^{c\times n\times hf}$. Optionally, if the resolution prefix tuning module is activated (as explained later), the resolution prefix is concatenated with $\bm{X}_{h}$. For notational simplicity, we denote the concatenated tensor with $\bm{X}_{h}$ as well.
@@
The TTM backbone consists of $L$ levels, each comprising $M$ TTM blocks with identical patch configurations. Each TTM block is comprised of a patch partition model, a vanilla TSMixer block, and a patch merging block. A TTM block in the $i$-th level, $i=1,\ldots,L$, receives the processed data $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}$ from the earlier block, and performs the following operations on it. Patch Partition block at every level $i$ increases the number of patches by a factor of $K_{i}$ and reduces the patch dimension size by the same factor by reshaping $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}\text{to }\bm{X}_{h}^{i}\in\mathbb{R}^{c\times(n\cdot K_{i})\times(hf/K_{i})}$, where $K_{i}=2^{(L-i)}$. Note that, we set $hf=2^{m}$ for some integer $m\geq L$. Then, TSMixer is applied to the adapted data $\bm{X}_{h}^{i}$. Finally, the output from TSMixer is again reshaped to its original shape (i.e., $\mathbb{R}^{c\times n\times hf}$) in the patch merging block. Note that, as we go deeper into the network, the number of patches decreases while the patch dimension size increases leading to adaptive patching which helps in better generalization as we pretrain with multiple datasets together. This idea of adaptive patching is popular and very successful in the vision domain (E.g. Swin transformers Liu et al. (2021) and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pretraining with diverse TS datasets. Figure 1(b) shows the TTM backbone for $L=3$ and $M=2$.
@@
Data Augmentation via Downsampling: A significant challenge in TS pretraining datasets is the scarcity of public datasets at specific resolutions. To overcome this, we employ a downsampling technique for high-resolution datasets, generating multiple datasets at lower resolutions. For example, from a one-second resolution dataset, we derive datasets at minute and hour resolutions. Importantly, the original high-resolution dataset remains within the pool of pretraining datasets. This methodology significantly augments the number of datasets for each resolution which greatly improves the model performance (Section 4.5).
@@
Resolution Prefix Tuning: This technique explicitly learns and incorporates a new patch embedding as a prefix into the input data based on the input resolution type (see Figure 1(b)). Similar to the concept of prefix tuning Li and Liang (2021), this approach provides an explicit signal to the model about the resolution type for resolution-conditioned modeling. First, we map every resolution to a unique integer, which is then passed through an embedding layer to project it to the hidden dimension, $hf$. Subsequently, we expand the embedding across all channels to have a representation of shape $c\times 1\times hf$. This module is optional for the TTM backbone, particularly beneficial when the context length ($sl$) is short. In those scenarios, automatically detecting the resolution becomes a challenge for the model. Hence, by explicitly fusing the resolution information as a prefix, we can enhance the model’s ability to learn effectively across resolutions.
@@ 3.2 Fine-tuning Workflow
In the fine-tuning workflow, we deal with data from the target domain that has no overlap with the pretraining datasets. We have three options here: (a) In Zero-shot forecasting, we directly use the pretrained model to evaluate on the test part of the target data, (b) In Few-shot forecasting, we utilize only a tiny portion (5-10%) of the train part of the target data to quickly update the pretrained weights of the decoder and head, and subsequently, evaluate it on the test part, (c) In Full-shot forecasting, we finetune the pretrained weights of the decoder and head on the entire train part of the target data, and then, evaluate on the test part.
@@
The backbone is completely frozen during fine-tuning, and still operates in a channel-independent univariate fashion. However, the TTM decoder can be fine-tuned via channel-mixing (for multivariate) or a channel-independent (for univariate) way based on the nature of the target data. If pure multivariate modeling is needed, then the channel-mixer block in all the TSMixer components (see Figure 1(b)) in the decoder gets enabled to explicitly capture the channel correlation between the channels. The forecast head and reverse normalization perform similar operations as in the pretraining stage. The fine-tuning also optimizes the forecasting objective with MSE loss. This thoughtful multi-level design choice ensures our backbone excels in channel-independent pretraining, enabling effective temporal correlation modeling across diverse datasets. Simultaneously, the decoder handles target-data-specific tasks like channel-correlation modeling and fine-tuning. In addition, if the target data has exogenous variables, then an exogenous mixer block is applied to the actual forecasts as explained next.
@@
Exogenous Mixer Block: As described in Section 2, the future values of the exogenous channels are known in advance. Let the forecast from the forecast head be $\hat{\bm{Y}}\in\mathbb{R}^{c\times fl}$. Let the channels $\bm{x}_{0},\cdots,\bm{x}_{N_{1}}$ denote the target variables and $\bm{x}_{N_{1}+1},\cdots,\bm{x}_{N_{1}+N_{2}}$ denote all exogenous variables with their future values known. First, we replace the forecast values for the exogenous channels with the true future values ($\bm{Y}$) and transpose it: $\hat{\bm{Y}}_{e}=\left[\hat{\bm{y}}_{0},\cdots,\hat{\bm{y}}_{N_{1}},\bm{y}_{N_{1}+1},\cdots,\bm{y}_{N_{1}+N_{2}}\right]\in\mathbb{R}^{fl\times c}$. Next, to learn inter-channel lagged correlations, we patch $\hat{\bm{Y}}_{e}$ into a series of overlapped windows (i.e., patching with stride$=1$) to create a new tensor: $\hat{\bm{Y}}_{e,p}\in\mathbb{R}^{fl\times\Delta\times c}$, where $\Delta=2\cdot l+1$ with $l$ being the context length to incorporate on either side of a time point [^4] This needs padding $\hat{\bm{Y}}_{e}$ with zeros of length $l$ on both sides.. Subsequently, we pass $\hat{\bm{Y}}_{e,p}$ through a vanilla TSMixer block with channel mixing enabled. Thus, the lagged dependency of the forecasts for the target channels on the exogenous channels is seamlessly learned. Finally, we attach a linear head to produce the forecasts for the target channels which is then reshaped as $\hat{\bm{Y}}\in\mathbb{R}^{N_{1}\times fl}$. Figure 1(c) depicts this procedure.

### 17 Jan 2024 v3

# Tiny Time Mixers (TTMs): Fast Pre-trained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

@@ 1 Introduction
Related Work: Recent advances in multivariate forecasting have been marked by the advent of transformer-based Vaswani et al. (2017) approaches, exemplified by models like PatchTST Nie et al. (2023), Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and FEDFormer Zhou et al. (2022). These models have demonstrated notable improvements over traditional statistical and ML methods. Furthermore, architectures based on MLP-Mixer Tolstikhin et al. (2021), such as TSMixer Ekambaram et al. (2023), have emerged as efficient transformer alternatives, boasting 2-3X reduced compute and memory requirements with no accuracy compromise compared to their transformer counterparts. However, none of these advanced approaches have successfully demonstrated the ability to create general pre-trained models that can successfully transfer the learning to unseen target TS dataset, in a similar way as popularly witnessed in NLP and vision tasks. This is very challenging in the TS domain due to the diverse nature of the datasets across applications and the limited public availability of TS data for pre-training. There are existing self-supervised pre-training TS approaches using masked modeling and contrastive learning techniques such as SimMTM Dong et al. (2023) and TF-C Zhang et al. (2022) that offer transfer learning between two datasets when carefully selected based on the dataset properties. However, they fail to provide universal transfer learning capabilities across datasets. Consequently, there has been a recent growing trend to employ pre-trained large language models (LLMs) for TS forecasting, treating it as a cross-domain transfer learning task. These universal cross-transfer approaches, specifically recent works such as LLMTime Gruver et al. (2023) and GPT4TS Zhou et al. (2023) yield promising results in few/zero-shot forecasting approaches. These models are bootstrapped from GPT-2/3 or LLAMA-2 with suitable tokenization strategies to adapt to time-series domains.
@@
However, these LLM based TS approaches do not explicitly handle channel correlations and exogenous support in the context of multivariate forecasting. Moreover, these large models, with billions of parameters, demand significant computational resources and runtime. Hence, in this paper, we focus on building pre-trained models from scratch solely using TS data. Unlike language, which has abundant public pre-training data in terabytes, time-series data is relatively scarce, very diverse and publicly limited. Its scarcity leads to overfitting when pre-training “large” models solely on time-series data. This prompts a question: Can smaller models pre-trained purely on limited public diverse TS datasets give better zero/few-shot forecasting accuracy? Surprisingly, the answer is yes! Toward this, we propose Multi-level Tiny Time Mixers (TTM), a significantly smaller model ($\leq$1M parameters) based on the lightweight TSMixer architecture, exclusively trained on diverse TS corpora for effective zero/few-shot multivariate TS forecasting via transfer learning.
@@
In particular, TTM is pre-trained using multiple public datasets ($\sim$ 244M samples) from the Monash data repository [^2] Accessible at https://forecastingdata.org/ Godahewa et al. (2021)). Note that the datasets exhibit considerable diversity in terms of characteristics, such as the different domains, temporal resolution [^3] Resolution refers to the sampling rate of the input time series (e.g., hourly, 10 minutes, 15 minutes, etc.) (spanning from second to daily), lengths, and number of channels. Pre-training on such heterogeneous datasets cannot be handled directly by TSMixer or existing state-of-the-art (SOTA) models. Hence, TTM proposes the following enhancements to the TSMixer architecture: (i) Adaptive Patching across layers, considering the varied suitability of patch lengths for different datasets, (ii) Dataset Augmentation via Downsampling to increase coverage and samples across different resolutions, (iii) Resolution Prefix Tuning to explicitly embed resolution information in the first patch, facilitating resolution-conditioned modeling, particularly beneficial in scenarios with short history lengths. Moreover, our approach leverages multi-level modeling, where TTMs are first pre-trained in a channel-independent way and then seamlessly integrate channel mixing during fine-tuning to model target data-specific channel-correlations and exogenous infusion
@@
  • Amidst the prevalence of large pre-trained models demanding significant compute and training time (in weeks), our work is the first to showcase the efficacy of building Fast and Tiny Pre-trained models ($\leq$1M parameters) exclusively trained on Public TS datasets in a flash of just few hours (4-8 hours, 6 A100 GPUs). TTM successfully demonstrates transfer learning to diverse, unseen target datasets for zero/few-shot forecasting, addressing the data scarcity issues prevalent in time series.
  • Pre-training on heterogeneous multi-resolution datasets cannot be handled effectively by TSMixer or other SOTA models. Hence, we propose various architectural and training enhancements, such as adaptive patching, data augmentation via downsampling, and (an optional) resolution prefix tuning for robust pre-training.
@@ 2 TTM Components
Let $\bm{X}\in\mathbb{R}^{c\times sl}$ be a multivariate time series of length $sl$ and $c$ number of channels. The forecasting task can be formally defined as predicting the future values $\bm{Y}\in\mathbb{R}^{c^{\prime}\times fl}$ given the history $\bm{X}$. Here, $fl$ denotes the forecast horizon and $c^{\prime}$ denotes number of forecast channels, where $c^{\prime}\leq c$. The predictions from the model are denoted by $\hat{\bm{Y}}\in\mathbb{R}^{c^{\prime}\times fl}$. In a general multivariate forecasting task, each channel or variable falls into one of the following categories: (a) Target variables (mandatory):corresponding to the channels for which forecasts are required, (b) Exogenous variables (optional):encompassing (i) uncontrolled variables that may influence the forecasts and assumed to be known or estimated for the forecast period (e.g. weather), and (ii) control variables whose future values during the forecast horizon can be manipulated to govern the behavior of the target variables. (e.g. discount in sales forecasting, operator controls in industrial applications). In TTM, uncontrolled and control variables are treated similarly, as both are considered available during forecasting.
@@ 3 TTM Workflows
TTM works in 2 stages: pre-train and fine-tune (Figure 1(a)).
3.1 Pre-training Workflow
@@
In the pre-training stage, we train the model on a large collection of public datasets from the Monash data repositoryGodahewa et al. (2021). Since the primary focus of TTM is forecasting, pre-training is modeled with a direct forecasting objective. TTM is first pre-trained in a univariate fashion with independent channels on all the existing datasets. Due to varied channel counts in pre-training datasets, modeling multivariate correlations is not feasible here; it is addressed later during fine-tuning. Multivariate pre-training datasets are initially transformed into independent univariate time series $(\bm{X}_{1},\cdots,\bm{X}_{N})\in\mathbb{R}^{c(=1)\times sl}$. These are pre-processed (Section 2.2), and subsequently fed into the TTM backbone for multi-resolution pre-training. The output of the backbone $\bm{X}_{h}^{L}\in\mathbb{R}^{(c=1)\times n\times hf}$ is passed through the decoder and forecast head to produce the forecast $\hat{\bm{Y}}\in\mathbb{R}^{(c=1)\times fl}$ which is then reverse-normalized to bring back to the original scale. We pre-train the TTM with mean squared error (MSE) loss function calculated over the forecast horizon: $\mathcal{L}=||\bm{Y}-\hat{\bm{Y}}||_{2}^{2}$. Thus for a given input context length $sl$ and forecast length $fl$, we get a pre-trained model capturing the common temporal forecasting dynamics and seasonal patterns as observed in the overall pre-training data.
3.1.1 Multi-Resolution Pre-training via TTM Backbone
@@
The majority of the pre-training happens in the TTM backbone. The primary challenge with the proposed pre-training technique is that the pre-training data is diverse and has multiple resolutions. There are two main options for pre-training: conducting separate pre-training for each resolution type or pre-training using all resolution data collectively. While it’s common to train a model per resolution type to overcome challenges in learning diverse seasonal patterns, this leads to diminished training data for each resolution due to limited data availability. Consequently, this motivated the exploration of pre-training a single model using datasets from all resolutions. To achieve this, we propose the following 3 enhancements.
@@
Adaptive Patching: The TTM backbone is crafted with an adaptive patching architecture where different layers of the backbone operate at varying patch lengths and numbers of patches. Since each dataset in the pre-training corpora may perform optimally at a specific patch length, this approach greatly aids in generalization when diverse datasets with different resolutions are introduced. As shown in Figure 1(b), the patched data $\bm{X}_{p}\in\mathbb{R}^{c\times n\times pl}$ is passed through a embedding layer to project it to the patch hidden dimension, $\bm{X}_{h}\in\mathbb{R}^{c\times n\times hf}$. Optionally, if the resolution prefix tuning module is activated (as explained later), the resolution prefix is concatenated with $\bm{X}_{h}$. For notational simplicity, we denote the concatenated tensor with $\bm{X}_{h}$ as well.
@@
The TTM backbone consists of $L$ levels, each comprising $M$ TTM blocks with identical patch configurations. A TTM block in the $i$-th level, $i=1,\ldots,L$, receives the processed data $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}$ from the earlier block. Each TTM block is further comprised of a patch partition model, a vanilla TSMixer block, and a patch merging block. Patch Partition block at every level $i$ increases the number of patches by a factor of $K_{i}$ and reduces the patch dimension size by the same factor by reshaping $\bm{X}_{h}^{(i-1)}\in\mathbb{R}^{c\times n\times hf}\text{to }\bm{X}_{h}^{i}\in\mathbb{R}^{c\times(n\cdot K_{i})\times(hf/K_{i})}$, where $K_{i}=2^{(L-i)}$. Note that, we set $hf=2^{m}$ for some integer $m\geq L$. Then, TSMixer is applied to the adapted data $\bm{X}_{h}^{i}$. Finally, the output from TSMixer is again reshaped to its original shape (i.e., $\mathbb{R}^{c\times n\times hf}$) in the patch merging block. Note that, as we go deeper into the network, the number of patches decreases while the patch dimension size increases leading to adaptive patching which helps in better generalization as we pre-train with multiple datasets together. This idea of adaptive patching is popular and very successful in the vision domain (E.g. Swin transformers Liu et al. (2021) and we are the first to port it successfully to the time-series domain to resolve multi-resolution issues in pre-training with diverse TS datasets. Figure 1(b) shows the TTM backbone for $L=3$ and $M=2$. Please note that adaptive patching is enabled only in the backbone and not in the decoder, which is designed to be very lightweight.
@@
Data Augmentation via Downsampling: A significant challenge in TS pre-training datasets is the scarcity of public datasets at specific resolutions. To overcome this, we employ a downsampling technique for high-resolution datasets, generating multiple datasets at lower resolutions. For example, from a one-second resolution dataset, we derive datasets at minute and hour resolutions. Note that, the original high-resolution dataset remains within the pool of pre-training datasets. This methodology significantly augments the number of datasets for each resolution which greatly improves the model performance (Section 4.5).
@@
Resolution Prefix Tuning: This technique explicitly learns and incorporates a new patch embedding as a prefix into the input data based on the input resolution type (see Figure 1(b)). Similar to the concept of prefix tuning Li and Liang (2021), this approach provides an explicit signal to the model about the resolution type for resolution-conditioned modeling. First, we map every resolution to a unique integer, which is then passed through an embedding layer to project it to the hidden dimension, $hf$. Subsequently, we expand the embedding across all channels to have a representation of shape $c\times 1\times hf$. This module is optional for the TTM backbone, particularly beneficial when the context length ($sl$) is short. In these scenarios, automatically detecting the resolution becomes a challenge for the model. Hence, by explicitly fusing the resolution information as a prefix, we can enhance the model’s ability to learn effectively across resolutions.
@@ 3.2 Fine-tuning Workflow
In the fine-tuning workflow, we deal with data from the target domain that has no overlap with the pre-training datasets. We have three options here: (a) In Zero-shot forecasting, we directly use the pre-trained model to evaluate on the test part of the target data, (b) In Few-shot forecasting, we utilize only a tiny portion (5-10%) of the train part of the target data to quickly update the pre-trained weights of the decoder and head, and subsequently, evaluate it on the test part, (c) In Full-shot forecasting, we fine-tune the pre-trained weights of the decoder and head on the entire train part of the target data, and then, evaluate on the test part.
@@
The backbone is completely frozen during fine-tuning, and still operates in a channel-independent univariate fashion. However, the TTM decoder can be fine-tuned via channel-mixing (for multivariate) or a channel-independent (for univariate) way based on the nature of the target data. If pure multivariate modeling is needed, then the channel-mixer block in all the TSMixer components (see Figure 1(b)) in the decoder gets enabled to explicitly capture the channel correlation between the channels. The forecast head and reverse normalization perform similar operations as in the pre-training stage. The fine-tuning also optimizes the forecasting objective with MSE loss. This thoughtful multi-level design choice ensures our backbone excels in channel-independent pre-training, enabling effective temporal correlation modeling across diverse datasets. Simultaneously, the decoder handles target-data-specific tasks like channel-correlation modeling and fine-tuning. In addition, if the target data has exogenous variables, then an exogenous mixer block is applied to the actual forecasts as explained next.
@@
Exogenous Mixer Block: As described in Section 2, the future values of the exogenous channels are known in advance. Let the forecast from the forecast head be $\hat{\bm{Y}}\in\mathbb{R}^{c\times fl}$. Let the channels $\bm{x}_{0},\cdots,\bm{x}_{c^{\prime}}$ denote the target variables and $\bm{x}_{c^{\prime}+1},\cdots,\bm{x}_{c}$ denote all exogenous variables with their future values known. First, we replace the forecast values for the exogenous channels with the true future values ($\bm{Y}$) and transpose it: $\hat{\bm{Y}}_{e}=\left[\hat{\bm{y}}_{0},\cdots,\hat{\bm{y}}_{c^{\prime}},\bm{y}_{c^{\prime}+1},\cdots,\bm{y}_{c}\right]\in\mathbb{R}^{fl\times c}$. Next, to learn inter-channel lagged correlations, we patch $\hat{\bm{Y}}_{e}$ into a series of overlapped windows (i.e., patching with stride$=1$) to create a new tensor: $\hat{\bm{Y}}_{e,p}\in\mathbb{R}^{fl\times\Delta\times c}$, where $\Delta=2\cdot l+1$ with $l$ being the context length to incorporate on either side of a time point [^4] This needs padding $\hat{\bm{Y}}_{e}$ with zeros of length $l$ on both sides.. Subsequently, we pass $\hat{\bm{Y}}_{e,p}$ through a vanilla TSMixer block with channel mixing enabled. Thus, the lagged dependency of the forecasts for the target channels on the exogenous channels is seamlessly learned. Finally, we attach a linear head to produce the forecasts for the target channels which is then reshaped as $\hat{\bm{Y}}\in\mathbb{R}^{c^{\prime}\times fl}$. Figure 1(c) depicts this procedure.